35,000 dollars were made off of 2,000 tickets sold. Student tickets were 10$ & non student tickets were 20$. How many students bought tickets & how many non student bought tickets?
Answer by TimothyLamb(4379)

s + n = 2000
n = 2000 - s
10s + 20n = 35000
---
10s + 20n = 35000
10s + 20(2000 - s) = 35000
10s + 40000 - 20s = 35000
10s = 40000 - 35000
s = 5000/10
---
s = 500
n = 1500
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php